OpenStudy (anonymous):

Polar coordinates of a point are given. Find the rectangular coordinates of the point. (2.1, 2pie/9)

OpenStudy (anonymous):

I think it is pie that confuses me

OpenStudy (anonymous):

\(x=2.1\cos(\frac{2\pi}{9})\) and \(y=2.1\sin(\frac{2\pi}{9})\) and a calculator for this one i think because i have no idea what that is

OpenStudy (anonymous):

x=2.099844112 y=.0255872301

OpenStudy (anonymous):

that is not what i get http://www.wolframalpha.com/input/?i=2.1+cos%282pi%2F9%29

OpenStudy (anonymous):

i think i don't have my calculator set right

OpenStudy (anonymous):

{1.60869, 0.642788}

OpenStudy (anonymous):

better be in "radian" mode

OpenStudy (anonymous):

so x= 1.6 and y= 1.3

OpenStudy (anonymous):

no it is in degree that is the problem

OpenStudy (anonymous):

thank you